what number must be added to the numerator and subtracted from the denominator of the fraction 17/32 to make a fraction equal to 3/4.
Answer by stanbon(75887)

what number must be added to the numerator and subtracted from the denominator of the fraction 17/32 to make a fraction equal to 3/4.
====
Let the number be "x"::
(17+x)/(32-x) = 3/4
-----
4(17+x) = 3(32-x)
68 + 4x = 96 - 3x
7x = 28
x = 4
-----------
Cheers,
Stan H.
-------------